OpenStudy (anonymous):

find the inverse (if it exists) of f(x) 5x-4

jimthompson5910 (jim_thompson5910):

f(x) = 5x - 4 y = 5x - 4 ... replace f(x) with y x = 5y - 4 ... swap x and y now solve for y the solution you get will be the inverse

OpenStudy (anonymous):

y= 4x/5?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

\[\large x = 5y - 4\] \[\large x+4 = 5y\] \[\large 5y = x+4\] \[\large y = \frac{x+4}{5}\]

jimthompson5910 (jim_thompson5910):

so the inverse function is \[\large f^{-1}(x) = \frac{x+4}{5}\]

OpenStudy (anonymous):

oh i have to add the x to 4

OpenStudy (anonymous):

okay!

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

do you mind helping with one more?(:

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

what's your question

OpenStudy (anonymous):

for f(x)+ 4x+1 and g(x) = x^2 -2x-4 find (g-f)(x)

jimthompson5910 (jim_thompson5910):

typo in f(x)+ 4x+1 ???

OpenStudy (anonymous):

and then simplify it. this one seems hard

jimthompson5910 (jim_thompson5910):

(g-f)(x) means g(x) - f(x)

OpenStudy (anonymous):

yes f(x) =

jimthompson5910 (jim_thompson5910):

so you are subtracting the two functions...ie the two expressions that make up the functions

jimthompson5910 (jim_thompson5910):

so... (g-f)(x) = g(x) - f(x) (g-f)(x) = (x^2 -2x-4) - (4x+1) (g-f)(x) = x^2 -2x-4 - 4x-1 ... don't forget to distribute fully (g-f)(x) = ???

OpenStudy (anonymous):

x^2+2x-3

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

unfortunately it's not

OpenStudy (anonymous):

dang it!

jimthompson5910 (jim_thompson5910):

-2x - 4x = ??

jimthompson5910 (jim_thompson5910):

its ok

OpenStudy (anonymous):

-6x?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

how about -4 - 1

jimthompson5910 (jim_thompson5910):

-4 - 1 = ???

OpenStudy (anonymous):

-5

jimthompson5910 (jim_thompson5910):

(g-f)(x) = g(x) - f(x) (g-f)(x) = (x^2 -2x-4) - (4x+1) (g-f)(x) = x^2 -2x-4 - 4x-1 (g-f)(x) = x^2 - 6x - 5

jimthompson5910 (jim_thompson5910):

and that's all there is to it

OpenStudy (anonymous):

you are brilliant thank you

jimthompson5910 (jim_thompson5910):

thanks for the compliment and I'm glad to be of help